OpenStudy (anonymous):

Hey guys need help with a question. The question is 3 diffrent rectangulers have a area of 28 square units what are the possible whole numbers dimensions of the rectangles? PLZ HELP ME GUYS:D

OpenStudy (anonymous):

@RhondaSommer

OpenStudy (anonymous):

im so sorry i need to do this and type my essay but i forgot my notebook so im frustruated :(

RhondaSommer (rhondasommer):

well... 4*7 28*1 14*2

RhondaSommer (rhondasommer):

they all equal 28 when multiplied.

OpenStudy (anonymous):

so its 28?

RhondaSommer (rhondasommer):

no. its 4*7 = 28. thats one rectangle. 28*1 = 28 thats second rectangle 14*2 = 28 thats a third rectangle

RhondaSommer (rhondasommer):

those are your dimentions

OpenStudy (anonymous):

so the possible whole dimensions are 4,7,28,1,14,2

OpenStudy (anonymous):

nvm the confused part

RhondaSommer (rhondasommer):

yep. :)

OpenStudy (anonymous):

ok thx wait so if i have to write 16 as a power in 2 difrent ways how would it be?

OpenStudy (anonymous):

sorry dats another question

OpenStudy (anonymous):

@RhondaSommer

RhondaSommer (rhondasommer):

as a power? like x^2=16 or x*y=16

OpenStudy (anonymous):

it just says write 16 as a power 2 diffrent ways

OpenStudy (anonymous):

@atreyu6s3x6

OpenStudy (anonymous):

whats the problem here?

OpenStudy (anonymous):

ok its above

RhondaSommer (rhondasommer):

umm i think it means x*y=16 ... 4*4=16 and 16*116

RhondaSommer (rhondasommer):

16*1=16 sorry

OpenStudy (anonymous):

ok thx im so sorry but thx

RhondaSommer (rhondasommer):

only thing that makes sense really why are you sorry

OpenStudy (anonymous):

well lets say ive keep asking u questions

RhondaSommer (rhondasommer):

your fine XD i habe no life anyway

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

what grade are u in :D

OpenStudy (anonymous):

or do u prefer not to aswer that

RhondaSommer (rhondasommer):

10th but im graduating next year

RhondaSommer (rhondasommer):

:)

OpenStudy (anonymous):

well i say u will graduate :D

RhondaSommer (rhondasommer):

no like im graduating a year early

OpenStudy (anonymous):

well gtg do other question and type :D bye

OpenStudy (anonymous):

yea i get it :D

RhondaSommer (rhondasommer):

you can tag me if u need to...

OpenStudy (anonymous):

thx for ALLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLLL the help and trust me i will tag u :D